OpenStudy (anonymous):

What is the inverse of the given relation?y = 3x + 9

OpenStudy (babyslapmafro):

or in other words, set the equation equal to x(y) instead of y(x)

OpenStudy (anonymous):

@sirm3d

OpenStudy (sirm3d):

the idea of inverse is to solve the independent variable in terms of the dependent variable. as for this problem \[y=3x+9\] transpose 9, then divide by 3 \[\frac{ y-9 }{ 3 }=x\] as for the final answer, switch variables in the final equation.

OpenStudy (anonymous):

y = 3x - 3?

hero (hero):

You can switch variables first, then solve. That doesn't matter.

OpenStudy (anonymous):

o so y = 3x + 9

hero (hero):

x = 3y + 9

OpenStudy (anonymous):

if not i confused myself

OpenStudy (anonymous):

ooooo x first ok

OpenStudy (anonymous):

i dont have x = 3x + 9 as an anwer

OpenStudy (sirm3d):

@hero now she's confused by you.

hero (hero):

That's because I provided the first step. There are more steps of course. By the way, the first step is x = 3y + 9

OpenStudy (anonymous):

yep lol

hero (hero):

Maybe I should provide all the steps again.

OpenStudy (anonymous):

yes please lol

OpenStudy (sirm3d):

we're in the final step, and you intruded. that's nice.

OpenStudy (anonymous):

wait we were in the final step ugh

hero (hero):

Bro, I assure you, she was lost even before I said a word.

OpenStudy (anonymous):

honestly yes because i dont understand half of this i dont even know what f(x) is

OpenStudy (sirm3d):

nice assumption. she hasn't answer my lsat question.

hero (hero):

It wasn't an assumption. Some things I just know.

OpenStudy (anonymous):

@Hero can you list the steps please

hero (hero):

\[ \\ y = 3x + 9 \\ x = 3y + 9 \\ x - 9 = 3y \\\frac{x - 9}{3} = y \]

OpenStudy (anonymous):

i dont have that as an answer

hero (hero):

I suppose you have \[\frac{x - 9}{3} = f(x)\]

hero (hero):

Or maybe even \[\frac{x - 9}{3} = f^{-1}(x)\]

OpenStudy (babyslapmafro):

\[y=3x+9\] \[y-9=3x\] \[\frac{ y-9 }{ 3 }=x\]

hero (hero):

^incomplete @Babyslapmafro

OpenStudy (babyslapmafro):

why is that

hero (hero):

Because x still has to be the input of the function.